What volume of hydrogen at STP is released when 8.35 g Mg reacts? (Assume that the amount of HCl is enough for the reaction.)


1
Expert's answer
2020-10-26T14:45:10-0400

Mg+2HCl=MgCl2+H2

n(H2)=n(Mg)=8.35/24.3=0.344 mol

V(H2) = 0.334*22.4 = 7.70 l